At 30.0 K, a container has an unknown pressure. What is the new temperature if the pressure is decreased to 33% of the original?


1
Expert's answer
2022-05-11T14:31:03-0400

Given:

P1 = P

T1 = 30.0 K

V1 = V2 = constant

P2 = 0.33 P

T2 = unknown


Solution:

Since the volume and amount of gas remain unchanged, Gay-Lussac's law can be used.

Gay-Lussac's law can be expressed as: P1 / T1 = P2 / T2

P1T2 = P2T1

To find the new temperature, solve the equation for T2:

T2 = P2T1 / P1

T2 = (0.33 P × 30.0) / (P) = 9.9

T2 = 9.9 K


Answer: The new temperature is 9.9 K